please help me solve this problem and show how it is done:
cuberoot(x)-5=3

Answer by sudhanshu_kmr(1152)

cuberoot(x)-5=3
=> cuberoot(x) = 3 + 5
=> cuberoot(x)= 8 =
=> x = 8^3 = 512